OpenStudy (anonymous):

MEDAL AND FAN! PLEASE HELP! When Tyson opens and lays a cereal box out flat, he sees that the top and the bottom of the box both measure 11 inches by 8 inches, the sides of the box both measure 11 inches by 6 inches, and the front and back of the box both measure 8 inches by 6 inches. What is the surface area of Tyson's cereal box?

OpenStudy (gorv):

sum all these

OpenStudy (gorv):

top=11*8 bottom=11*8 side1=11*6 sid2=11*6 front=8*6 back=8*6 surface area= sum of area of all the faces =88+88+66+66+48+48

OpenStudy (gorv):

404

OpenStudy (gorv):

@Zejinida

OpenStudy (gorv):

check it out

OpenStudy (anonymous):

So when you times them all together you get the answer?

OpenStudy (anonymous):

So its 404?

OpenStudy (gorv):

yep time up all the faces

OpenStudy (gorv):

yeah it's 404

OpenStudy (anonymous):

Alrighty thank you :3